A body vibrating with simple harmonic motion has a frequency of 4hz and an a amplitude of 0.15m.Calculate the maximum values of the acceleration and velocity
The acceleration and velocity at a point 0.09 from the equilibrium position of motion
1
Expert's answer
2019-03-27T05:47:43-0400

If x = 0.15sin25.12t m, then velocity v = 0.15*25.12cos25.12t = 3.77cos25.12t m/s and acceleration a = -94.7sin25.12t m/s2. When x = 0.09 m sinωt = 0.6 therefore a1 = -56.82 m/s2 and velocity v1 = 3.77*0.8 = 3.02 m/s.

So, vm = 3.77 m/s, v1 = 3.02 m/s, am = 94.7 m/s2, a1 = -56.82 m/s2.